3x/x-2 = 3x-5/x -1
Answer by checkley79(3341)

3x/x-2 = 3x-5/x-1 cross multiply.
(x-2)(3x-5)=3x(x-1)
3x^2-6x-5x+10=3x^2-3x
3x^2-3x^2-11x+3x+10=0
-8x+10=0
-8x=-10
x=-10/-8
x= 5/4 ans.
Proof:
(3*5/4)/(5/4-2)=(3*5/4-5)/(5/4-1)
(15/4)/(5/4-8/4)=(15/4-20/4)/(5/4-4/4)
(15/4)/(-3/4)=(-5/4)/(1/4)
(15/4)(-4/3)=-(5/4)(4/1)
-60/12=-20/4
-5=-5